# Koopman Operators in Robot Learning

### v1

@@ Abstract
Koopman operator theory offers a rigorous treatment of dynamics and has been emerging as a powerful modeling and learning-based control method enabling significant advancements across various domains of robotics. Due to its ability to represent nonlinear dynamics as a linear operator, Koopman theory offers a fresh lens through which to understand and tackle the modeling and control of complex robotic systems. Moreover, it enables incremental updates and is computationally inexpensive making it particularly appealing for real-time applications and online active learning. This review comprehensively presents recent research results on advancing Koopman operator theory across diverse domains of robotics, encompassing aerial, legged, wheeled, underwater, soft, and manipulator robotics. Furthermore, it offers practical tutorials to help new users get started as well as a treatise of more advanced topics leading to an outlook on future directions and open research questions. Taken together, these provide insights into the potential evolution of Koopman theory as applied to the field of robotics.
@@ I Introduction
The emphasis in the last decade on deep learning has contributed to reliance on ‘big data’ in robotics [14, 15], driven by slow algorithms operating on offline data that have been labeled at scale (e.g., the massive availability of labeled images, contributing to the prioritization of vision as a sensing modality in robotics). In contrast, our effort herein centers on the following motivating question. If robots—with their physical bodies and embodied learning constraints—are going to function in novel settings without relying on offline data, what tools are available that are appropriate for runtime learning using only ‘small data?’
Figure 1: Left: Overview of the Koopman-based runtime learning, modeling, and control of robots. Right: Key components of the implementation.
Koopman operators provide a partial answer to this question. The main way they are integrated within core robotics foundational research (learning, modeling, and control) is summarized in Fig. 1. The essential ingredients in Koopman operators are an infinite basis of observables, a linear operator predicting the evolution of those observables, and some method for approximating the operator using finite calculations. These ingredients are sufficient to represent any nonlinear system, including nonsmooth and discontinuous ones [16]. These operators have only recently made their way into the field of robotics, both theoretically and experimentally. Machine learning (data-driven) techniques using Koopman operators are suited to robotics needs—both in terms of empirical properties and formal properties. Empirically, Koopman operators only require sparse data sets, making them amenable to runtime computation using only small datasets. Moreover, this enables complex adaptive behavior (e.g., spontaneous control response to unmodeled turbulent flow). Formal properties include applying physically relevant properties to otherwise unknown models (e.g., stability, invariance, or symmetry), the ability to directly calculate information measures for continuous active learning, the benefit of applying linear control techniques (e.g., LQR) to nonlinear systems and the ease of certifying stability (e.g., constructive Control Lyapunov Functions [17]), and guarantees on sample efficiency. These benefits collectively contribute data-driven techniques that can be implemented synchronously with robotic execution, rather than in an offline batch processing manner.
The emergence of Koopman-based models presents a potential paradigm shift, offering a unified framework for modeling nonlinear dynamics by providing more efficient and versatile representations of robotic behaviors and dynamics that have the following distinct advantages.
  1. 1.
    Interpretability: Unlike many machine learning approaches that represent the input-output relationship as a black box neural network (NN), the Koopman operator theory provides dynamical model descriptions of the underlying system rooted at principled geometric and algebraic properties that can be leveraged to explain the performance of data-driven approximations.
  2. 2.
    Data-efficiency: The Koopman operator theory demands only a limited number of measurements when compared to most of the NN-based methods, making it suitable for real-time implementation.
  3. 3.
    Linear Representation: The Koopman operator leads to a linear representation of the dynamics, enabling the use of linear systems tools and methods for analysis, estimation, and control synthesis.
As Koopman operator theory continues to rapidly evolve, several studies have delved into algorithmic enhancements and validation techniques for modeling and control purposes. While the main body of relevant research results started to appear only within the past five years, it is rapidly growing. A literature search with the specific keywords “Koopman” and “Robot*” yields over 17,800 results on Google Scholar.111 Data are current up to April 24th, 2024. The publications with the most relevance and their number of citations are shown in Fig. 2. The graphs clearly illustrate the increasing trend in this area. While existing review articles have extensively discussed data-driven methodologies [18] and operator-based algorithms [19], recent review articles on Koopman operator theory have predominantly focused on theoretical analyses [20, 21, 22], encompassing topics such as operator estimation, spectral properties analysis, and controller design [23]. More recently, there has been a surge in reviews specific to soft robotics [24]. Here, our goal is to offer a comprehensive introduction to the Koopman operator, complete with a tutorial featuring executable code for implementing different procedures, along with recent advancements in its application within the field of robotics. We aim to elucidate the benefits of Koopman-based approaches for tasks such as modeling, state estimation, and control. Additionally, we seek to consolidate and highlight recent trends in research efforts related to the practical utilization of Koopman operator theory across various robotic systems.
Figure 2: Literature review results of the Koopman operator theory in robotics with application instances.
The review is structured to be accessible to both newcomers and those with prior experience in Koopman and robotics. The recommended reading roadmap is outlined in Fig. 3. Specifically, in Sections II and III, we provide an overview of the Koopman operator along with a practical tutorial, augmented by code available online.222 The code is available on Github. This serves as a quick reference for those new to the field, allowing them to grasp the basic concepts of how Koopman operator theory functions and how to implement it for robot modeling and control. Readers can choose to stop here and experiment with the provided code with their systems. Then, the interested reader may proceed to Section VI to explore state-of-the-art applications across various robotic platforms, such as aerial robots, ground robots, robot arms, underwater robots, soft robots, and others, including examples in multi-agent systems. For readers seeking more in-depth understanding, Section IV goes into theoretical details of the Koopman operator, covering aspects like Koopman operators in discrete-time and continuous-time systems, Koopman eigenfunctions as well as Koopman invariant spaces, data-driven approximation approaches, and strategies for handling systems with inputs. Continuing to Section V, we present the implementation details of the Koopman operator in robotic systems. This encompasses guidance on data collection, selection of lifting functions, and controller design. In the concluding Section VII of the review, we provide a summary of the covered materials and discuss several open challenges and potential future directions in the field.
Figure 3: The illustration of recommended reading sequence. For readers who are new to the Koopman operator in robotics, we recommend following the green line to read this review paper. In addition, interested readers who seek to learn more about this area are suggested to follow the orange line.
II The Koopman Operator at A Glance
II-A The Koopman Operator
Let $x\in\mathcal{X}\subseteq\mathbb{R}^{N_{x}}$ be the state vector for a discrete-time dynamical system. The system’s propagation rule is represented by the nonlinear function $T$ (Fig. 4). Through the process of lifting the original states $x$ using the observable functions $x\mapsto g(x)$, the system’s dynamics can be redefined in a new space where a linear operator, denoted as the Koopman operator $\mathcal{K}$, describes the dynamics. In other words, although $T$ and $\mathcal{K}$ act on different spaces, they encapsulate the same dynamics. For example, given the current state $x$, one can propagate it to the subsequent time step and observe it through two pathways: either by employing $T$ to compute $T(x)$ and observe the resultant state (the bottom route) or by utilizing an observable function, applying $\mathcal{K}$ and evaluating it at $x$ (the top route). The concept of the “equivalence” or “substitution” offers several advantages. (1) It enables a global linear representation of the nonlinear dynamics $T$, thus enabling the application of techniques designed for linear systems. (2) It facilitates the estimation of the underlying dynamics by learning the linear operator in real time, thus eliminating the need for least-square regression of the nonlinear function, which in general requires a large amount of data.
II-B Data-driven Prediction via Koopman Theory
Consider a discrete-time system [^3] One can also describe similar constructions for continuous-time dynamical systems, but we delay its presentation to Section IV. expressed as
$x_{t+1}=T(x_{t})\enspace.$ (1)
Consider a vector space $\mathcal{F}$ of complex-valued functions (known as observables) with the system’s state space as their domain. One can think of the observables $g\in\mathcal{F}$ as mapped or lifted states. The evolution of the observables $g$ using the Koopman operator $\mathcal{K}:\mathcal{F}\to\mathcal{F}$ associated with system (1) is
$\mathcal{K}g=g\circ T,\quad\forall g\in\mathcal{F}\;,$
where $\circ$ denotes function composition. To ensure $\mathcal{K}$ is well defined, $\mathcal{F}$ must be closed under composition with $T$. This condition might force $\mathcal{F}$ to be infinite-dimensional if one insists that $\mathcal{F}$ contains specific predefined functions (such as functions that return the full state values).
The Koopman operator propagates the system one step forward in time according to
$\mathcal{K}g(x_{t})=g\circ T(x_{t})=g(T(x_{t}))=g(x_{t+1})\;.$ (2)
The lifted states $g(x)$ essentially transform the original (nonlinear) system into a linear one evolving at the Koopman space which can afford the design and implementation of a variety of linear controllers (such as LQR).
@@ II-B Data-driven Prediction via Koopman Theory
Figure 4: A glance at the Koopman Operator Theory.
To avoid dealing with infinite-dimensional spaces, a finite-dimensional representation of the Koopman operator can be employed instead. Thus, finite-dimensional subspaces of $\mathcal{F}$ that are invariant under the Koopman operator play a crucial role. Formally, let $\mathcal{S}\subset\mathcal{F}$ be a finite-dimensional Koopman-invariant subspace and restrict the action of the Koopman operator to $\mathcal{S}$ as $\mathcal{K}\!\!\restriction_{\mathcal{S}}:\mathcal{S}\to\mathcal{S}$. Since $\mathcal{S}$ is finite-dimensional, given any basis for it, one can represent the operator $\mathcal{K}\!\!\restriction_{\mathcal{S}}$ by a matrix. Formally, let $\mathbf{\Psi}$ be a vector-valued function whose elements form a basis for $\mathcal{S}$, then there exists a matrix $K\in\mathbb{C}^{\dim(\mathcal{S})\times\dim(\mathcal{S})}$ such that
$\mathcal{K}\mathbf{\Psi}=\mathbf{\Psi}\circ T=K\mathbf{\Psi}\;,$ (3)
$\mathbf{\Psi}(x_{t+1})=K\mathbf{\Psi}(x_{t})\;.$ (4)
By defining $z_{t}:=\mathbf{\Psi}(x_{t})$, (4) can be written as $z_{t+1}=Kz_{t}$ which is in linear form. If the space $\mathcal{S}$ contains the state observables $g_{i}(x)=x^{i}$, where $x^{i}$ is the $i$th element of state $x$, then one can choose the basis $\mathbf{\Psi}$ in a way that contains $g_{i}$’s as its elements.444 This is referred to as the full-state observability assumption [25]. More generally, state observables can be written as a linear combination of elements of $\mathbf{\Psi}$, if they are contained in $\mathcal{S}$. In this case, the lifted linear system (4) captures the complete information of the original nonlinear system (1). Moreover, the eigendecomposition of matrix $K$ allows one to identify the Koopman eigenfunctions and eigenvalues. We refer the reader to Section IV for more details.
It is generally challenging (sometimes impossible) to find exact finite-dimensional invariant subspaces that contain all the state observables. A popular way to tackle this issue is via approximations. Suppose the space $\mathcal{F}$ is equipped with an inner product, given any arbitrary space $\mathcal{S}$ (not necessarily invariant under the Koopman operator). In that case, one can consider the operator $\mathcal{P}_{\mathcal{S}}\mathcal{K}:\mathcal{F}\to\mathcal{F}$, where $\mathcal{P}_{\mathcal{S}}$ is the orthogonal projection operator on $\mathcal{S}$. For this operator, the space $\mathcal{S}$ is invariant by construction. Therefore, we can restrict the action of this operator to $\mathcal{S}$ as $\mathcal{P}_{\mathcal{S}}\mathcal{K}\!\!\restriction_{\mathcal{S}}:\mathcal{S}\to\mathcal{S}$ and given a basis $\mathbf{\Psi}$ for $\mathcal{S}$, the operator admits a matrix representation
$\mathcal{P}_{\mathcal{S}}\mathcal{K}\mathbf{\Psi}=\hat{K}\mathbf{\Psi}\;,$
where $\hat{K}\in\mathbb{C}^{\dim(\mathcal{S})\times\dim(\mathcal{S})}$. In this case, one can write an approximate version of (4) as follows
$\mathbf{\Psi}(x_{t+1})=[\mathcal{K}\mathbf{\Psi}](x_{t})\approx[\mathcal{P}_{\mathcal{S}}\mathcal{K}\mathbf{\Psi}](x_{t})=\hat{K}\mathbf{\Psi}(x_{t})\;.$ (5)
By defining $z_{t}:=\mathbf{\Psi}(x_{t})$, the previous equation can be written in an approximate linear form $z_{t+1}\approx\hat{K}z_{t}$, enabling the use of efficient linear methods to approximate the system’s behavior.
In practice, we typically have access to discretely sampled measurement data of the system. This can then be used to obtain a finite-dimensional approximated matrix of the infinite-dimensional Koopman operator and these Koopman-based components. Of all the approximation approaches, one of the most popular algorithms is the Extended Dynamic Mode Decomposition (EDMD) approach [25]. In particular, if $\mathcal{F}$ is the $L_{2}$ space based on the empirical measure on the data set, the procedure above corresponds to EDMD, which allows one to compute the matrices $K$ and $\hat{K}$ solely based on data by solving least-squares problems (see Section IV-B for details).
The Koopman operator theory, along with its data-driven estimation approach EDMD, was initially formulated without accounting for inputs. Several ways have been proposed to include control inputs. For example, control inputs can be integrated into the definition of $\mathbf{\Psi}$ by treating them as an augmented state [26]. Then, the process described above can be employed similarly for estimating the system $x_{t+1}=T(x_{t},u_{t})$. A detailed discussion on how to deal with this case is provided in Section IV-C.
III Implementation At a Glance
This section summarizes how to train a Koopman-based model and use it for control synthesis. There are mainly three steps: data collection (from physical experiments or simulations), model training, and control synthesis as shown in Algorithm 1. Next, we present a high-level overview of each step; more details are also provided in Section V.
The quality of the training data greatly impacts the ability to learn an accurate representation of the underlying system and, subsequently, apply effective control. The aim is to collect sequences of state and control measurements that allow one to learn how an initial state and control input map to the next state: $(x_{t},u_{t})\mapsto x_{t+1}$. To learn a globally accurate representation, training data must sample the entire state and control spaces $\mathcal{X}$ and $\mathcal{U}$, respectively. While this is practically infeasible, one should strive to capture training data that cover as much of the operating space of the robot as possible. A non-uniform data distribution will bias the learned model toward the more frequently observed states in the training dataset.
Given a training dataset, one can learn a Koopman model after selecting the basis functions as well as the optimization criteria and parameters. The basis functions determine the structure of the Koopman representation (i.e. linear, bilinear, nonlinear) [27] and can be either composed based on a pre-determined dictionary of functions or learned (e.g., via a neural network). The optimization, too, can differ and, besides optimizing for modeling error, it can include regularization terms, constraints that enforce, or cost terms that promote model properties (i.e. stability). Modeling errors can be assessed by different metrics, and this choice affects the answer one obtains, as well as the computational complexity of solving the corresponding optimization. The most direct implementation involves a least-squares regression solution that can be computed efficiently in real-time. After training a Koopman model, and depending on the structure of the dynamics, one can deploy appropriate feedback control methods. For example, linear Koopman models are used with LQR or linear MPC, and bilinear Koopman models can pair up with nonlinear MPC.
Figure 5: Overall structure of the modeling and control of robotic systems with the Koopman operator theory.
Data Collection:
  • States and Inputs Determination: Based on the control objective and sensors available on the system, determine the state and input.
  • Data Collection: Collect data with approaches like those discussed in Section V-A.
  • Data Preprocessing (as required): Perform data curation: normalize when the magnitude among different states varies too much; filter and/or average to remove noise; interpolate between measurements to ensure consistent temporal spacing across all training samples, etc.
Model training (Online or Offline):
  • Lifting Function Selection: select lifting functions, e.g., polynomials or NN-based approaches, etc. Refer to Sec. V-B for ways to define the dictionary of observables.
  • Koopman operator estimation: Utilize $X$, $U$ to estimate the Koopman operator $K$.
  • Koopman eigenfunctions calculation (Optional): obtain the Koopman eigenfunction of the operator for analysis if needed.
  • Compute Koopman Mode, eigenvalue and eigenvector expressions (Optional)
Control Synthesis
  • Design the Controller: design control structure, e.g., LQR, based on the estimated Koopman model.
  • 1
    Feedback Signal Collection (if applied): obtain
real-time observations from the robot.
  • Obtain Control Signal: calculate the control input to drive the robotic system to achieve the control objective.
  • Algorithm 1 Implementation Algorithm
    An implementation overview of Koopman operator theory in robotic systems is included in the supplementary tutorial (https://shorturl.at/ouE59). We also recommend investigating the impact of different dictionary selections and experimenting with various approaches to estimate the Koopman operator using the code provided by [28]. Additionally, designing and implementing different controllers can be explored using examples provided in [29], while the scripts used in [30, 31] offer insights into active data collection.
    IV The Koopman Operator and Its Approximation
    We present the Koopman operator for both discrete- and continuous-time dynamical systems and present data-driven approaches to approximate it with finite-dimensional representations. Key notation is summarized in Table I.
@@ IV The Koopman Operator and Its Approximation
    $\{\mathcal{K}^{t}\}$ Koopman operator group of continuous-time system
    $K_{\operatorname{EDMD}}$ Approximated Koopman operator via EDMD
    $\mathcal{L}_{G}$ Koopman generator
@@
    $\mathbf{\Psi}$ Vector-valued observation
    IV-A Koopman Operator
    We briefly review the Koopman operator for discrete- and continuous-time systems following [32, 33].
    IV-A1 Discrete-Time Case

### v2

@@ Abstract
Koopman operator theory offers a rigorous treatment of dynamics and has been emerging as an alternative modeling and learning-based control method across various robotics sub-domains. Due to its ability to represent nonlinear dynamics as a linear (but higher-dimensional) operator, Koopman theory offers a fresh lens through which to understand and tackle the modeling and control of complex robotic systems. Moreover, it enables incremental updates and is computationally inexpensive, thus making it particularly appealing for real-time applications and online active learning. This review delves deeply into the foundations of Koopman operator theory and systematically builds a bridge from theoretical principles to practical robotic applications. We begin by explaining the mathematical underpinnings of the Koopman framework and discussing approximation approaches for incorporating inputs into Koopman-based modeling. Foundational considerations, such as data collection strategies as well as the design of lifting functions for effective system embedding, are also discussed. We then explore how Koopman-based models serve as a unifying tool for a range of robotics tasks, including model-based control, real-time state estimation, and motion planning. The review proceeds to a survey of cutting-edge research that demonstrates the versatility and growing impact of Koopman methods across diverse robotics sub-domains: from aerial and legged platforms to manipulators, soft-bodied systems, and multi-agent networks. A presentation of more advanced theoretical topics, necessary to push forward the overall framework, is included. Finally, we reflect on some key open challenges that remain and articulate future research directions that will shape the next phase of Koopman-inspired robotics. To support practical adoption, we provide a hands-on tutorial with executable code at https://shorturl.at/ouE59.
@@ I Introduction
The emphasis in the last decade on deep learning has contributed to reliance on ‘big data’ in robotics [14, 15], driven by slow algorithms operating on offline data that have been labeled at scale (e.g., the massive availability of labeled images, contributing to the prioritization of vision as a sensing modality in robotics). In contrast, our effort herein centers on the following motivating question. If robots—with their physical bodies and embodied learning constraints—are going to function in novel settings without highly relying on offline data, what tools are available that are appropriate for runtime learning using only ‘small data?’
@@
Figure 1: Overview of the utility and integration of Koopman operator theory and practice in robotics.
Koopman operators provide a partial answer to this question. The main way they are integrated within core robotics foundational research (modeling, control and estimation, and motion planning) is summarized in Fig. 1. The essential ingredients in Koopman operators are an infinite basis of observables, a linear operator predicting the evolution of those observables, and some method for approximating the operator using finite calculations. These ingredients are sufficient to represent any nonlinear system, including nonsmooth and discontinuous ones [16]. These operators have only recently made their way into the field of robotics, both theoretically and experimentally. Machine learning (data-driven) techniques using Koopman operators are suited to robotics needs—both in terms of empirical properties and formal properties. Empirically, Koopman operators only require sparse data sets, making them amenable to runtime computation using only small datasets [17]. Owing to their runtime computation affordances, Koopman operators can facilitate adaptive system behavior (e.g., spontaneous control response to unmodeled turbulent flow), that is hard to achieve otherwise. Formal properties include applying physically relevant properties to otherwise unknown models (e.g., stability, invariance, or symmetry), the ability to directly calculate information measures for continuous active learning, the benefit of applying linear control techniques (e.g., LQR) to nonlinear systems and the ease of certifying stability (e.g., constructive Control Lyapunov Functions [18]), and guarantees on sample efficiency. These benefits collectively contribute data-driven techniques that can be implemented synchronously with robotic execution, rather than in an offline batch processing manner.
Koopman operators can offer a unified framework for modeling nonlinear dynamics by providing more efficient and versatile representations of robot behavior, which in turn can facilitate reactive motion planning and adaptive control. Such models have the following three distinct advantages. (1) Interpretability: Unlike many machine learning approaches that represent the input-output relationship as a black box neural network (NN), the Koopman operator theory provides dynamical model descriptions of the underlying system rooted at principled geometric and algebraic properties that can be leveraged to explain the performance of data-driven approximations. (2) Data-efficiency: The Koopman operator theory demands only a limited number of measurements when compared to most of the NN-based methods, making it suitable for real-time implementation. (3) Linear Representation: The Koopman operator leads to a linear representation of the dynamics, enabling the use of linear systems tools and methods for analysis, estimation, and control synthesis.
Koopman operator theory applied to robotics is relatively new (the bulk of works discussed herein have appeared within the past five years) and continues to rapidly evolve to introduce algorithmic enhancements and validation techniques for modeling and ‘downstream’ tasks such as controller design (Fig. 1). While existing review articles have extensively discussed data-driven methodologies [19] and operator-based algorithms [20], recent review articles on Koopman operator theory have predominantly focused on theoretical analyses [21, 22, 23], encompassing topics such as operator estimation, spectral properties analysis, and controller design [24]. Further, a recent review focused exclusively on soft robotics [25].
Here, our goal is to offer a comprehensive introduction to the Koopman operator and how it applies to robotics with a range of specific examples across various robotics sub-domains. We seek to highlight the benefits of Koopman-based approaches for modeling, control and estimation, and motion planning. Additionally, we seek to consolidate and highlight recent trends in research efforts related to the practical utilization of Koopman operator theory across various robotic systems. To this end, the review is organized as follows. We begin in Section II with the fundamentals of Koopman operator theory, where we introduce the core mathematical concepts, data-driven estimation, and prediction methods. Section III transitions into the use of Koopman-based models in robotics and their usage for control, state estimation, and planning. Section IV presents a system-level breakdown of Koopman applications across robotics sub-domains (including manipulation, legged locomotion, soft and continuum robots, aerial and underwater vehicles, and multi-agent systems). To strengthen the ties to control theory and summarize the important theoretical underpinnings that are required to push forward the state of the underlying theory, Section V presents more advanced topics on Koopman-based modeling and control for general (nonlinear) systems. Lastly, Section VI outlines future research directions and Section VII concludes the review.
II Fundamentals of Koopman Operator Theory
In 1931, Bernard O. Koopman [26] demonstrated that a nonlinear dynamical system can be represented using an infinite-dimensional linear operator acting on a Hilbert space of measurement functions of the system’s state. Owing to this, the Koopman operator has recently gained attention in robotics for various tasks, such as modeling, control, planning, and observer design. In this section, we introduce the fundamentals of Koopman Operator Theory. We begin by summarizing the core mathematical concepts,111 A more extensive and rigorous treatise is provided in Section V. and continue to present how to estimate the operator directly from data. Since the Koopman operator was originally formulated for unforced systems, we also explain how Koopman approximations can be extended to systems with inputs. Key notation is summarized in Table I.
@@ II Fundamentals of Koopman Operator Theory
$K$ Approximated finite-dimensional Koopman operator
@@
$\mathbf{\Psi}$ Vector-valued observable function
II-A Koopman Operator Fundamentals